### v1

# Spatial Policy: Guiding Visuomotor Robotic Manipulation with Spatial-Aware Modeling and Reasoning

Abstract
Vision-centric hierarchical embodied models have demonstrated strong potential for long-horizon robotic control. However, existing methods lack spatial awareness capabilities, limiting their effectiveness in bridging visual plans to actionable control in complex environments. To address this problem, we propose Spatial Policy (SP), a unified spatial-aware visuomotor robotic manipulation framework via explicit spatial modeling and reasoning. Specifically, we first design a spatial-conditioned embodied video generation module to model spatially guided predictions through a spatial plan table. Then, we propose a spatial-based action prediction module to infer executable actions with coordination. Finally, we propose a spatial reasoning feedback policy to refine the spatial plan table via dual-stage replanning. Extensive experiments show that SP significantly outperforms state-of-the-art baselines, achieving a 33.0% average improvement over the best baseline. With an 86.7% average success rate across 11 diverse tasks, SP substantially enhances the practicality of embodied models for robotic control applications. Code and checkpoints are maintained at https://plantpotatoonmoon.github.io/SpatialPolicy/.
Introduction
In recent years, hierarchical embodied models combining high-level planning and low-level action have achieved remarkable success in long-horizon robotic manipulation (Brohan et al. 2024; Cheang et al. 2024; Ze et al. 2024; Li et al. 2025b). Among these approaches, vision-based methods, known as visuomotor robotic manipulation, have emerged as a particularly promising direction. By leveraging rich visual cues, these methods enable effective cross-hierarchical information transfer across diverse robotic embodiments and environments (Du et al. 2023; Tian et al. 2024; Ko et al. 2023; Soni et al. 2025; Black et al. 2023; Li et al. 2025a). Specicially, visuomotor robotic manipulation employs a generative model to predict trajectories of future video, which are subsequently translated into executable actions through a action prediction module.
Despite significant advances, existing visuomotor robotic manipulation frameworks lack spatial awareness ability, failing to effectively bridge visual plans to executable actions in complex environments (Tian et al. 2024; Black et al. 2023). For instance, AVDC (Ko et al. 2023) lacks explicit spatial modeling in its video generation process, resulting in physically implausible predictions, such as robotic arms passing through walls. Feedback module of CLOVER (Bu et al. 2024) enhances dynamic environment adaptation, but its inability to process spatial relationships prevents handling of occluded target scenarios (Tian et al. 2024).
To tackle this problem, we investigate the spatial-aware visuomotor robotic manipulation via explicit spatial modeling and reasoning, which introduces three key challenges: (1) how to generate videos that explicitly model spatial relations, such as object-goal positioning and task alignment; (2) how to utilize the generated videos to guide action prediction in a way that preserves spatial consistency and task relevance; and (3) how to reason over spatial cues during execution to enable real-time feedback that effectively bridges visual and control.
To address these challenges, we propose Spatial Policy (SP), a unified framework for robotic manipulation that integrates spatial-aware video generation, action prediction, and feedback. The core idea is to unify spatial modeling and reasoning via a spatial plan table that guides video generation, enables spatially consistent action prediction, and supports feedback-driven refinement during execution. Specifically, (1) we propose a Spatial-Conditioned Embodied Video Generation module that generates visual predictions guided by a structured spatial plan table, which encodes object-goal relations and directional cues to ensure spatial coherence in imagined futures, addressing the challenge of modeling spatially meaningful video trajectories; (2) we introduce a Spatial-Based Action Prediction module that infers executable actions by capturing spatial motion dynamics between predicted frames using spatial coordinates, tackling the challenge of translating visual predictions into spatially consistent actions; (3) we develop a Spatial Reasoning Feedback Policy that refines the spatial plan and video execution through dual-stage replanning based on VLM feedback and rule-based corrections, solving the challenge of real-time spatial reasoning during policy execution. Extensive experiments demonstrate that our model outperforms existing baselines, achieving 86.7% performance. The contributions of this paper are summarized as follows:
  • Spatial policy for vision-action alignment. We propose SP, a spatially grounded visuomotor policy framework that explicitly models and reasons about spatial structures to align visual prediction with executable actions.
  • Modular design with structured spatial modeling and reasoning. SP consists of three synergistic modules: a spatial-conditioned video generation module guided by a structured spatial plan table, a spatial-based action prediction that preserves spatial consistency, and a spatial reasoning feedback policy that refines execution through real-time spatial replanning.
  • Comprehensive evaluation and strong performance. We conduct extensive experiments on the MetaWorld benchmark to assess the impact of spatial integration, including ablations on video generation, action prediction, and feedback mechanisms. SP outperforms prior methods by over 30%, demonstrating the effectiveness of spatial reasoning in visuomotor robotic manipulation.
Related Work
Video Generation for Embodied Manipulation
Video generation has recently been explored as a visual planning mechanism for embodied agents. AVDC (Ko et al. 2023) generates action-annotated video plans from raw RGB inputs using dense correspondencesenabling policy learning without action supervision. However, due to the lack of explicit spatial reasoning in both generation and action prediction, limiting precision in manipulation tasks. VideoAgent (Soni et al. 2025) refines video plans via self-conditioning to reduce hallucinations and improve visual quality. However, it inherits the lack of explicit spatial reasoning from AVDC in both generation and action prediction, limiting precision in geometry-sensitive manipulation tasks. Clover (Bu et al. 2024) introduces a closed-loop framework combining video diffusion planning and feedback control via an embedding-based error signal, achieving strong results in long-horizon tasks. However, its spatial modeling is limited to depth-based cues and lacks explicit 3D spatial planning or structured geometric reasoning, which restricts generalization across diverse spatial layout. SEER (Tian et al. 2024)proposes a Predictive Inverse Dynamics Model (PIDM) that forecasts future visual states and infers actions end-to-end, closing the loop between vision and control. While it achieves strong performance through large-scale training, it lacks explicit spatial reasoning modules and does not leverage structured spatial plans, which may limit interpretability and adaptability in geometry-sensitive tasks. SuSIE (Black et al. 2023) leverages an image-editing diffusion model to generate visual subgoals for a low-level goal-conditioned policy, enabling better generalization in unstructured environments. However, its subgoal generation relies on appearance-based edits without explicit incorporation of 3D spatial relationships, limiting its precision in spatially complex manipulation scenarios.
Spatial Reasoning in Vision-Language Embodied Systems
Spatial understanding has been addressed in embodied AI from several perspectives. SpatialVLM (Chen et al. 2024a) and GraspCorrect (Lee, Hong, and Kim 2025) leverage pretrained vision-language models to extract chain-of-thought spatial cues and determine object relations in manipulation tasks. 3D-VLA (Zhen et al. 2024) explicitly encodes spatial semantics by aligning 3D vision with textual instructions. SPA (Zhu et al. 2025) utilizes point cloud reconstruction for spatial alignment in manipulation and navigation tasks. Meanwhile, LLM-Planner (Song et al. 2023)and related methods employ large language models to infer spatially structured plans in symbolic form.
Despite these advances, such methods either focus on semantic spatial understanding or operate in symbolic domains, lacking visual feedback and executable rollout. Moreover, most rely on handcrafted prompts or static spatial priors and do not support continuous video generation. In contrast, our work closes this gap by embedding structured spatial plans—automatically derived from environment geometry—into a generative video prediction framework, thereby tightly coupling spatial reasoning and visual synthesis for action-aware planning.
Spatial Conditioning in Generative Video Models
Recent advances in text-to-video diffusion models, such as AnimateDiff (Guo et al. 2024) and VideoCrafter2 (Chen et al. 2024b), achieve high-fidelity generation but lack controllability over agent behavior or scene dynamics. ScenarioDiff (Zhang et al. 2025) and DirecT2V (Hong et al. 2024) propose spatially guided generation using manually specified constraints like bounding boxes or motion paths. These constraints are often synthetic and decoupled from real embodied agent states, limiting their adaptability and semantic alignment with embodied control tasks.
Spatial Policy
We propose a visuomotor robotic manipulation framework that integrates structured spatial reasoning into the perception-to-control pipeline, enabling robots to generate spatially consistent visual plans and execute them reliably. As illustrated in Fig. 1, our system comprises three interconnected modules. First, the Spatial-Conditioned Embodied Video Generation module generates future video trajectories conditioned on a structured spatial plan table composed of atomic actions, directional vectors, and relative distances. Second, the Spatial-Based Action Prediction module predicts actions aligned with the generated video using the spatial coordinates. Third, the Spatial Reasoning Feedback Policy monitors spatial alignment during execution and dynamically refines both the spatial plan table and the visual predictions. Together, these modules form a closed-loop framework that maintains spatial consistency across video imagination, action prediction, and feedback-based reasoning.
Figure 1: Framework Overview of Spatial Policy. Our system comprises three modules: (1) Spatial-Conditioned Embodied Video Generation, which uses a Spatial Plan Table obtained via VLM reasoning over the spatial offset between the robot and the object, to guide diffusion model in generating spatially coherent video prediction. (2) Spatial-Based Action Prediction, which converts the generated video into executable actions using spatial coordinates. (3) Spatial Reasoning Feedback Policy, which enables real-time correction via a dual-stage replanning strategy, combining VLM-based video judgement and policy diagnostics to refine the spatial plan table for closed-loop control.
Spatial Conditioned Embodied Video Generation
To generate spatially grounded visual trajectories for visuomotor robotic manipulation, we adopt a conditional diffusion framework for future video prediction. The model takes as input an initial observation frame and a textual task description, and generates a sequence of future frames representing high-level visual plans. To enhance spatial awareness, we additionally introduce spatial plan table derived from spatial relationships as an auxiliary global condition.
Spatial Plan Table
To capture precise spatial intent, we extract the coordinates of the robot end-effector $\mathbf{p}_{\text{ee}}\in\mathbb{R}^{3}$ and the target object $\mathbf{p}_{\text{obj}}\in\mathbb{R}^{3}$ from the simulator, and compute their relative offset as $\Delta\mathbf{p}=\mathbf{p}_{\text{obj}}-\mathbf{p}_{\text{ee}}$. This spatial offset is passed into a pretrained visual-language model (GPT-4o) to produce a structured spatial plan. The generated plan consists of sequential subgoals, each encoding an atomic action with an action type, a direction vector, and a distance scalar, as illustrated in Fig. 1.
Encoding and Integration
Each subgoal is independently encoded into a fixed-dimensional latent vector. Concretely, the action type is mapped to an embedding $\mathbf{e}_{\text{act}}\in\mathbb{R}^{d}$ via a learnable embedding table, the direction vector is discretized and embedded into $\mathbf{e}_{\text{dir}}\in\mathbb{R}^{d}$ using another learnable embedding table, and the distance scalar is projected into $\mathbf{e}_{\text{dis}}\in\mathbb{R}^{d}$ through a learnable MLP. The combined subgoal embedding is given by:
$\mathbf{e}_{\text{subplan}}=\text{MLP}\left([\mathbf{e}_{\text{act}};\,\mathbf{e}_{\text{dir}};\,\mathbf{e}_{\text{dis}}]\right),$ (1)
All subgoal embeddings $\{\mathbf{e}_{1},\dots,\mathbf{e}_{n}\}$ are concatenated with the task text embedding $\mathbf{z}_{\text{task}}\in\mathbb{R}^{d}$ obtained via CLIP (Radford et al. 2021) to form the composite global condition:
$\mathcal{T}=\left[\mathbf{z}_{\text{task}};\,\mathbf{e}_{1};\,\dots;\,\mathbf{e}_{n}\right].$ (2)
This global condition is injected into all layers of the u-net through conditional normalization mechanisms, guiding the diffusion model to produce spatially grounded, instruction-aligned video sequences.
Conditional Diffusion Formulation
We follow the design of a u-net-based denoising diffusion probabilistic model, adapted from the Imagen architecture (Ho et al. 2022), consistent with AVDC (Ko et al. 2023). Given an initial frame $\mathbf{I}_{0}\in\mathbb{R}^{H\times W\times 3}$ and a composite condition vector $\mathcal{T}$ that incorporates both task instructions and structured spatial information, the model predicts a sequence of $T=7$ future frames $\mathbf{I}_{1:T}$, resulting in an 8-frame trajectory. The generative objective is to approximate the posterior distribution:
$p(\mathbf{I}_{1:T}\mid\mathbf{I}_{0},\mathcal{T}),$ (3)
where the training target is to denoise a Gaussian-perturbed version of the future frames. Specifically, the model minimizes the following mean squared error (MSE) loss:
$\mathcal{L}_{\text{MSE}}=\mathbb{E}_{\epsilon,t}\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{1-\beta_{t}}\cdot\mathbf{I}_{1:T}+\sqrt{\beta_{t}}\cdot\epsilon,\,t\mid\mathbf{I}_{0},\mathcal{T}\right)\right\|^{2},$ (4)
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is sampled noise, $\beta_{t}$ is the diffusion schedule, and $\epsilon_{\theta}$ is the model’s denoising network. Further training details are provided in the Appendix.
Spatial Based Action Prediction
To convert spatially grounded visual plans into executable actions, we design an Spatial Coordinate-Guided Action Prediction module built upon the diffusion policy framework (Chi et al. 2023; Ji et al. 2025), due to its strong capacity for modeling multimodal distributions and complex temporal dependencies, making it particularly well-suited for capturing the spatial variability and uncertainty inherent in visual planning. The model predicts multi-step actions that guide the robot from the current observation toward the visual goal implicitly defined by the generated video sequence, leveraging spatial coordinate to capture fine-grained spatiotemporal dynamics.
Multimodal Spatial Input Encoding
We condition our policy on three types of information: (1) the current image $\mathbf{I}_{\text{cur}}$, (2) the goal image $\mathbf{I}_{\text{goal}}$ (sampled from the generated visual plan within the next $k$ steps, i.e., $\mathbf{I}_{\text{goal}}=\mathbf{I}_{t+\delta}$ where $1\leq\delta\leq k$), and (3) the spatial coordinate of robotic arms $\mathbf{p}_{\text{cur}}\in\mathbb{R}^{3}$. These inputs capture both spatial and temporal motion cues. All inputs are encoded into compact feature vectors: the image pair is stacked along the channel dimension and passed through a shared ResNet encoder, while the spatial coordinate is encoded via a shallow MLP. The resulting modality-specific features are then concatenated to form a unified condition vector, which guides the denoising network in our diffusion policy-based action prediction model.
Diffusion Policy-based Action Prediction Model
We formulate action prediction as a conditional generative process. Specifically, we adopt a unet-based diffusion policy architecture to model the conditional distribution over action trajectories:
$p_{\theta}(\mathbf{a}_{0:H-1}\mid\mathbf{I}_{\text{cur}},\mathbf{I}_{\text{goal}},\mathbf{p}_{\text{cur}})$ (5)
where $\mathbf{a}_{0:H-1}$ denotes a sequence of $H$ actions, each including a 3-DoF end-effector pose and gripper control. We train the policy via denoising score matching, minimizing:
$\mathcal{L}_{\text{MSE}}=\mathbb{E}_{\mathbf{a},\epsilon,t}\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{1-\beta_{t}}\cdot\mathbf{a}+\sqrt{\beta_{t}}\cdot\epsilon,t\mid\mathbf{c}\right)\right\|^{2}$ (6)
where $\mathbf{c}$ denotes the concatenated global condition vector from all encoded inputs, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is Gaussian noise, and $t$ is the diffusion timestep. $\epsilon_{\theta}$ is the learned denoising network. At inference time, the model predicts the entire action trajectory in a denoising pass, and the resulting sequence is executed step-by-step in the environment. Further training details are provided in the Appendix.
Spatial Reasoning Feedback Policy
To ensure spatial consistency throughout execution, we introduce a Spatial Reasoning Feedback Policy that enables closed-loop correction by dynamically refining the Spatial Plan Table based on perceptual feedback. This module integrates a dual-stage feedback mechanism, operating at both the video validation and action execution levels.
During visual planning, we incorporate a vision-language model (VLM)-based consistency checker inspired by VideoAgent (Soni et al. 2025). A pretrained GPT-4o model is prompted to assess the spatial plausibility of the generated video—verifying object-goal alignment, manipulator trajectories, and overall spatial coherence. Only videos that pass this check are used for downstream control; otherwise, regeneration is triggered under the same spatial condition.
During action execution, we employ rule-based policy diagnostics to monitor temporal progress and detect failures such as positional drift or behavioral stagnation. When triggered, the system computes an updated spatial offset between the end-effector and the target object based on the latest observation. This offset is re-encoded into a refined spatial subplan via GPT-4o, forming an updated condition to replan both the video and action policy.
Through this real-time refinement of spatial intent and control guidance, our system maintains alignment between visual imagination and robotic execution, enabling robust long-horizon manipulation.
Experiments
Method door-open door-close basketball shelf-place btn-press btn-press-top
AVDC 30.7% 28.0% 21.3% 8.0% 34.7% 17.3%
AVDC-R 72.0% 89.3% 37.3% 18.7% 60.0% 24.0%
VA 40.0% 29.3% 13.3% 9.3% 38.7% 18.7%
VA-OR 82.7% 97.3% 40.0% 26.7% 73.3% 44.0%
SP 98.6% 100.0% 66.7% 80.0% 94.7% 96.0%
SP-R 98.6% 100.0% 73.3% 92.0% 96.0% 98.6%
Method faucet-close faucet-open handle-press hammer assembly Overall
AVDC 12.0% 17.3% 41.3% 0.0% 5.3% 19.6%
AVDC-R 53.3% 24.0% 81.3% 8.0% 6.7% 43.1%
VA 46.7% 12.0% 36.0% 0.0% 1.3% 22.3%
VA-OR 74.7% 46.7% 86.7% 8.0% 10.7% 53.7%
SP 93.3% 89.3% 100.0% 74.7% 24.0% 83.4%
SP-R 92.0% 96.0% 100.0% 73.3% 32.0% 86.7%
Table 1: Comparison of task success rates (%) across 11 MetaWorld tasks. Each method is evaluated over 800 episodes per task. R denotes variants equipped with a replanning mechanism. OR indicates methods that combine online reinforcement learning with replanning. Bold indicates best.
Datasets and Experimental Setups
Datasets and Environments. We follow the evaluation setup of (Ko et al. 2023) and conduct experiments in the Meta-World (Yu et al. 2020). Meta-World is a simulated benchmark designed for assessing robotic manipulation capabilities, consisting of 11 representative single-arm tasks (such as pushing, pulling, grasping, and inserting) performed by a Sawyer robotic arm. We run 25 random seeds across 11 tasks and 3 camera views, yielding over 800 rollouts in total, and evaluate 5 ablation studies in our model.
Baselines and Our Methods
  • AVDC: A two-stage control planning framework is employed, wherein a pretrained diffusion-based generative model synthesizes a video sequence, followed by the estimation of inter-frame motion using an optical flow network. The resulting motion is then used to inversely infer the corresponding action sequence.
  • AVDC-Replan: An extension of AVDC that incorporates a replanning mechanism. During execution, if action failure or robotic arm stagnation is detected, a new video is regenerated and a revised action plan is inferred.
  • VideoAgent: This method employs a self-supervised consistency mechanism to iteratively refine the generated video. Multiple candidate video plans are scored using GPT-4, and the most promising one is selected for execution.
  • VideoAgent-online-Replan: An extension of VideoAgent that continuously collects successful execution trajectories during deployment. These are used to update both the video generation and optimization models online. A replanning strategy is also introduced to regenerate videos when execution failures occur.
  • SP (Ours): We extract structured planning information from spatial context using VLM, and inject it into the diffusion model via global condition to effectively guide video generation. The generated video is subsequently translated into executable actions using a goal-conditioned policy.
  • SP-Replan (Ours): An extended version of our method that performs iterative video generation in the planning phase. During execution, if robotic arm stagnation is detected, a new video is regenerated and the action sequence is re-inferred.
Main Results
As shown in Table 1, our method achieves a success rate of 83.4% without requiring execution-time replanning, already outperforming all prior baselines. In contrast, methods such as AVDC and VA rely on task text to generate videos or pretrained optical flow to compute displacement coordinates for action prediction. These baselines are fragile during deployment, once a deviation or unexpected event occurs, they must fall back on explicit replanning to recover; otherwise, the robot stalls. By contrast, our approach leverages spatial plans to generate structured, goal-aligned videos and predicts executable actions that inherently adapt to spatial variations, leading to more robust and consistent task completion. SP-R further boosts performance to 86.7%, demonstrating the effectiveness of incorporating real-time spatial reasoning into the control loop.
On the subset of challenging tasks ( Assembly, Hammer, Faucet Open, Button Press Top, Shelf Place, and Basketball ), our method achieves an average success rate of 77.5%, significantly outperforming the best-performing baseline, which only reaches 29.2% . This corresponds to a relative improvement of 165.4%. On simpler tasks, our approach achieves near-perfect performance, often approaching 100% success.
Figure 2: Comparative visualization of spatial planning and video generation results for two representative tasks: Shelf Place and Basketball. The yellow line depicts the motion trajectory; the symbol marks the initial position, the symbol indicates the final task position, and the symbols denote intermediate steps along the path.
In the Shelf Place task, as shown in Figure 2, the baseline method generates an incorrect and inflexible trajectory (highlighted in the yellow box), resulting in task failure due to its inability to adapt to environmental variations. In contrast, our approach, enhanced with spatial understanding, enables the robotic arm to follow a dynamically planned path and successfully complete the placement task. Similarly, in the Basketball task (also illustrated in Figure 2), our method exhibits refined spatial reasoning by allocating more frames (highlighted in the green box) around the critical manipulation event—grasping the ball—leading to smoother and more precise motions. The baseline method, lacking this focused spatial adaptation, often misses essential interaction moments. These results underscore the benefits of integrating spatial awareness with dynamic feedback for robust and effective long-horizon robotic planning.
Figure 3: Visualization of the replanning-based spatial planning process for a manipulation task.
Figure 3 further demonstrates how our method leverages spatial feedback and iterative replanning to recover from failures and achieve successful task execution. Unlike baseline planning, which can fail when the initial path is invalid, our approach dynamically adapts, resulting in more robust and reliable performance.
Ablation Study
Ablations on the Video Generation Model. We compare three strategies in the video generation module to evaluate their impact on video quality and downstream control performance. (1) The baseline uses AVDC’s video generation approach, where a diffusion model is conditioned only on task text and the initial frame. (2) The Distance variant directly encodes the spatial offset between the robot and target object (spatial coordinates) and concatenates it with the task prompt as input. (3) The Subplan variant leverages a vision-language model to transform spatial coordinates into structured subplans that encode action type, direction, and distance, which are then injected into the diffusion model.
As shown in Table 2, both spatial variants outperform the AVDC baseline, demonstrating the utility of explicitly modeling spatial structures for enhanced visuomotor performance. Moreover, the subplan-based approach achieves further gains by abstracting raw spatial cues into structured guidance, highlighting the importance of structured spatial abstraction in facilitating precise and actionable visual plan generation.
Ablations on the Action Prediction Model. AVDC uses a pre-trained flow model to estimate pixel-wise motion between the current and goal frames, then converts the flow into a spatial target coordinate. The robot executes actions using a classical controller to move toward the point and applies gripper control based on heuristics. In contrast, our method is built on diffusion policy and directly predicts executable actions from the current frame, goal frame, and robot spatial coordinates. This design captures spatial and temporal context jointly and allows flexible adaptation to unexpected events such as object shifts or failures. Results in Table 3 show that our approach achieves higher success rates and better robustness, confirming that learning-based spatial reasoning provides more effective action prediction than coordinate-based control.
Ablation on Spatial Condition Injection Strategy in Video Generation. To investigate how spatial conditions affect video generation, we conduct an ablation study comparing two injection strategies for integrating the spatial plan table into the diffusion model. (1) The spatial plan table is concatenated with the task embedding and propagated uniformly across all layers of the U-Net backbone in the Global Cond variant directly. This approach provides consistent spatial guidance throughout the generation process. (2) Inspired by GLIGEN (Li et al. 2023), the spatial plan is injected only into intermediate U-Net layers via a cross-attention mechanism in the Local Cond variant, allowing localized spatial modulation while preserving the task-level prompt as a global condition.
Both strategies are trained under identical configurations and evaluated on MetaWorld. As shown in Table 4, Local Conditioning slightly improves the average task success rate and shows better performance on harder tasks (e.g., +5.3% on Hard). These results suggest that injecting spatial structure via localized attention can enhance spatial precision during video synthesis, thereby improving policy success rates, especially in more challenging environments.
Easy Medium Hard Very Hard Overall
AVDC 75.0% 48.7% 26.7% 48.0% 63.4%
Distance 93.5% 56.7% 25.3% 81.3% 79.4%
Subplan 96.8% 63.3% 24.0% 78.7% 82.4%
Table 2: Comparison of task success rates (%) using different video generation strategies across four levels of task difficulty. AVDC uses text and initial frame only; Distance adds raw spatial offsets; Subplan injects structured spatial subplans.
Easy Medium Hard Very Hard Overall
Flow 25.9% 10.7% 5.3% 8.0% 19.6%
Ours 75.0% 48.7% 26.7% 48.0% 63.4%
Table 3: Comparison of task success rates (%) between our method and AVDC’s action prediction module across four levels of task difficulty. Flow denotes the flow model-based action prediction used in AVDC; Ours denotes our spatial-based action prediction module.
Easy Medium Hard Very Hard Overall
Global Cond 96.8% 63.3% 24.0% 78.7% 82.4%
Local Cond 99.0% 62.0% 29.3% 80.0% 84.3%
Table 4: Comparison of task success rates (%) using different spatial condition injection strategies for video generation across four levels of task difficulty. Global Cond injects the spatial plan table across all U-Net blocks; Local Cond follows GLIGEN-style partial injection via cross-attention at intermediate layers.
Ablations on the Replanning Mechanism. We compare three settings to evaluate the effect of the replanning mechanism: (1) no replanning at all, (2) replanning only during the video generation stage, and (3) replanning during both the video generation and real execution stages. As shown in Table 5, task success improves when replanning is introduced.
Our experiments demonstrate that enabling replanning during both the imagination and real execution stages significantly enhances task success by facilitating dynamic adaptation to the evolving environment. While replanning solely during the imagination stage offers limited improvements, the integration of a vision-language-based evaluation module allows the system to selectively regenerate and retain higher-quality trajectories, effectively filtering out suboptimal predictions despite the inherent stochasticity of video generation. Without any replanning, the model’s performance deteriorates due to its inability to recover from unforeseen variations. Therefore, a combined replanning strategy is crucial for achieving robust and effective long-horizon robotic control.
Easy Medium Hard Very Hard Overall
NR 96.8% 63.3% 24.0% 78.7% 82.4%
GR 96.0% 70.7% 24.0% 80.0% 83.4%
GR&ER 97.3% 73.3% 32.0% 92.0% 86.7%
Table 5: Comparison of task success rates (%) using different replanning mechanisms across four levels of task difficulty. NR denotes no replanning; GR denotes generation replanning; ER denotes execution replanning. Bold indicates best.
Ablations on Finetuning with Additional Execution Data. To assess the impact of trajectory diversity on policy robustness, we compare two training data configurations for our model. (1) The model is trained exclusively on demonstrations where tasks are completed successfully from start to finish without failure. (2) The model is further finetuned on an expanded dataset that includes recovery trajectories where execution initially failed but eventually succeeded when resumed from intermediate states.
Finetuning on these recovery trajectories yields noticeable gains, particularly on challenging tasks. On btn-press-top, the success rate improves from 97.3% to 99.0%. More strikingly, on the more complex hammer task, performance increases from 68.0% to 78.7%. These results indicate that including partial-to-successful executions enhances the model’s robustness by exposing it to diverse recovery scenarios. Such data enables better generalization to states not frequently seen in the original dataset of consistently successful trials.
Hyperparameter Analysis
Effect of Diffusion Noise Seed in Video Generation. To assess the robustness of our video generation module under different stochastic conditions, we conduct experiments using our proposed method (SP) on four challenging MetaWorld tasks: basketball, shelf-place, hammer, and assembly. For each task, we sample six different diffusion noise seeds and report the task success rates. As shown in Figure 4, our approach consistently achieves high performance across seeds, with an average success rate around 77% and low variance. This indicates that the structured spatial subplans effectively anchor the generation process, ensuring the production of reliable and executable visual trajectories even under varying random conditions.
Figure 4: Task success rates on four challenging MetaWorld tasks under different diffusion noise seeds using our SP method, demonstrating consistent performance across seeds.
Effect of Hyperparameter in Action Prediction. We analyze key design choices in the action prediction module. First, we study the effect of goal frame selection range. Setting the goal frame as a random future frame within the next $k=20$ timesteps yields over 10% higher average success rates compared to restricting it to a narrower window between 17–20 timesteps ahead, indicating that allowing a broader temporal horizon improves generalization and task adaptability. Second, regressing the noise $\epsilon$ achieves better performance than directly predicting the denoised action $\hat{a}_{0}$, which contrasts with some prior findings (Chi et al. 2023). Lastly, while direct MLP-based action regression converges faster, the diffusion-based denoising model demonstrates superior performance and robustness across tasks.
Conclusion
We present SP, a spatial-aware visuomotor framework that integrates explicit spatial modeling and reasoning into robot manipulation. Central to our approach is a spatial plan table that encodes atomic actions, directions, and distances, enabling structured spatial guidance throughout the pipeline. Our polciy generates spatially coherent visual trajectories by conditioning a video diffusion model on spatial plan table, , which are then used to guide action prediction based on temporally aligned spatial coordinates. A dual-stage replanning mechanism enhances spatial reasoning by refining the spatial plan table and regenerating video and actions upon inconsistency, ensuring closed-loop correction. Experiments on MetaWorld show that SP improves success rates by over 33% compared to prior methods, validating the effectiveness of spatial modeling and reasoning.
Appendix A Inference Pipeline
As shown in alg. 1, Our inference process consists of three stages: spatial-conditioned embodied video generation, spatial-based action prediction, and spatial reasoning feedback policy. Given an initial observation and the current spatial coordinates of the robot and object, we compute a relative spatial offset $\Delta\mathbf{p}$, which is used to generate a spatial plan table via a VLM. The spatial plan table is fused with the task prompt and fed into a diffusion video generator to produce a goal-directed visual trajectory.
To infer executable actions, the action prediction model takes the predicted video sequence and encodes the current spatial state, represented solely by the robot end effector coordinate $\mathbf{p}_{\text{ee}}$, through a learnable MLP. This enables precise grounding of actions in the spatial context. During execution, a feedback module monitors task progress. If stagnation is detected, a two stage replanning mechanism is triggered: the spatial plan table is refined, and a new video and action sequence are generated to resume progress. This pipeline enables real time, spatially consistent decision making under dynamic conditions.
Input: Environment $\mathcal{E}$, Task prompt $\mathcal{T}_{0}$, Policy $\pi$
Output: Success flag $\mathcal{S}$
$\mathbf{I}_{0}\leftarrow\mathcal{E}.\texttt{Render}(),\;(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}})\leftarrow\mathcal{E}.\texttt{GetSpatialState}()$ ;
$\mathcal{S}_{0}\leftarrow\texttt{GPT4o}(\Delta\mathbf{p}=\mathbf{p}_{\text{obj}}-\mathbf{p}_{\text{ee}}),\;\mathcal{T}\leftarrow[\mathcal{T}_{0};\mathcal{S}_{0}]$ ;
repeat
 $\mathbf{V}_{1:T}\leftarrow\texttt{DiffusionGenerate}(\mathbf{I}_{0},\mathcal{T})$
until $\texttt{VLMCheck}(\mathbf{V}_{1:T})$ succeeds;
$\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\quad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
for $t=1$ to $H$ do
 $\mathcal{E}.\texttt{Step}(\mathcal{A}_{t})$ ;
 if IsStuck($\mathcal{E}$) then
    $\mathbf{I}_{t}\leftarrow\mathcal{E}.\texttt{Render}(),\;(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}})\leftarrow\mathcal{E}.\texttt{GetSpatialState}(),\;\mathcal{S}_{t}\leftarrow\texttt{GPT4o}(\Delta\mathbf{p}),\;\mathcal{T}_{t}\leftarrow[\mathcal{T}_{0};\mathcal{S}_{t}]$ ;
    repeat
       $\mathbf{V}_{1:T}\leftarrow\texttt{DiffusionGenerate}(\mathbf{I}_{t},\mathcal{T}_{t})$
    until $\texttt{VLMCheck}(\mathbf{V}_{1:T})$ succeeds;
    $\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\quad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
$\mathcal{S}\leftarrow\mathcal{E}.\texttt{CheckSuccess}()$ ;
return $\mathcal{S}$
Algorithm 1 Inference Pipeline
Appendix B Training Details: Video Generation and Action Prediction
Video Generation
Training Settings.
All models are trained on $128\times 128$ RGB frames sampled as 8-frame clips. We use the AdamW optimizer (Loshchilov and Hutter 2019) with an initial learning rate of $1\times 10^{-4}$ and a cosine learning rate schedule. Training is conducted for 100,000 steps, unless otherwise specified. The standard batch size is 8 for training and 16 for validation, with gradient accumulation set to 1. Mixed precision training is employed using FP16 and automatic mixed precision (AMP). A linear diffusion noise schedule is used with $\beta_{1}=1\times 10^{-4}$ and $\beta_{T}=0.02$ over 1000 steps. Exponential Moving Average (EMA) with a decay factor of 0.999 is updated every 10 training steps. Conditional dropout with a probability of 0.1 is applied to all models except the Baseline. Validation uses 1 or 2 samples per batch depending on the script. Unless otherwise specified, all training datasets are collected and generated from our own single successful executions. All experiments are conducted on a single NVIDIA RTX 4090 GPU (80GB memory), with full-precision (FP32) training completed within 2 days.
Inference.
During inference, all models use DDIM sampling with classifier-free guidance. Models conditioned on plans demonstrate improved semantic consistency and object interaction fidelity.
Model Variants and Conditioning.
We investigate four variants of our conditional diffusion models, differing mainly in conditioning inputs and architectural enhancements. Key differences and training datasets are summarized in Table 6.
Version Description
Baseline (Global Conditioning on Task) Standard UNet conditioned on global CLIP text embedding, serving as a task-aware video prior.
Global Conditioning on Task & Distance Adds a 9-dimensional gripper-object distance vector encoder for enhanced spatial grounding.
Global Conditioning on Task & Subplan Incorporates structured multi-step action plans as global conditioning signals.
Global Conditioning on Task & Subplan (Fine-tuned) Initially trained on expert data for 10k steps, then finetuned on additional datasets for 2k steps.
Local Conditioning on Task & Subplan Introduces cross-attention to fuse global and local plan features for fine-grained temporal conditioning.
Table 6: Comparison of video generation model variants highlighting key architectural or training differences.
Action Prediction
Training Settings.
Unlike the original Diffusion Policy (Chi et al. 2023), which conditions on past observations and actions, our model relies solely on the current observation, future observation, and spatial context. Specifically, instead of using a standard 3-channel RGB input, we concatenate the current frame and a sampled goal frame to form a 6-channel input image of resolution $128\times 128$, along with spatial coordinates. Given horizon $H=4$, the model predicts $H$ future actions from a single observation step ($T=1$), conditioned on the current frame, spatial coordinates, and a goal frame sampled within the next $K=20$ frames. We adopt a ResNet-50 visual encoder (He et al. 2016) pretrained on ImageNet (Deng et al. 2009), which remains trainable during training. The spatial coordinates are encoded via a MLP, which is also trained end-to-end. The policy network is implemented as a conditional U-Net with diffusion denoising, using a cosine noise schedule ($\beta_{1}=1\times 10^{-4}$, $\beta_{T}=2\times 10^{-2}$) over 100 steps. The model is optimized using AdamW with learning rate $\eta=1\times 10^{-4}$, weight decay $1\times 10^{-6}$, and batch size $B=512$ for both training and validation. The learning rate follows a cosine annealing schedule with $500$ warm-up steps. Exponential Moving Average (EMA) is applied with a maximum decay rate of $0.9999$ to stabilize training. All experiments are conducted on a single NVIDIA RTX 4090 GPU (40GB memory), with full-precision (FP32) training completed within 3 days.
Inference.
During inference, we adopt DDIM sampling with classifier-free guidance to generate action sequences. Conditioning on spatial plans significantly enhances semantic alignment and spatial fidelity, leading to more coherent and goal-directed behaviors.
Appendix C Tasks in Meta-World
In the Meta-World benchmark (Yu et al. 2020), we evaluated the performance on 11 representative robotic manipulation tasks shown in Figure 5, which are designed to assess various manipulation skills and generalization abilities of reinforcement learning algorithms. These tasks are categorized by their difficulty levels as follows:
  • Very Hard: Shelf Place: The agent must grasp an object and place it onto the target shelf. This task is particularly challenging due to the precision required in both grasping and placing the object accurately on the shelf.
  • Hard: Assembly: The agent grasps the handle position and inserts the ring into the slot to complete assembly. This task requires precise manipulation and coordination of multiple parts.
  • Medium: Hammer: The agent grasps the hammer handle and strikes forward to push the nail in with the hammer head. This task involves force control and spatial accuracy. Basketball: The task involves grasping the basketball, lifting it up, moving forward above the hoop, and placing it onto the hoop to complete the task. It requires coordinated movement and spatial awareness.
  • Easy: Door Open: The agent is required to push into the door handle grip area and then pull to open the door. Door Close: The agent needs to push into the door handle grip area and then push to close the door. Button Press: The agent moves horizontally to the button and pushes it in. Button Press Topdown: The agent moves above the button and presses down vertically. Faucet Close: The agent moves to the left side of the faucet handle and turns it right to close. Faucet Open: The agent moves to the right side of the faucet handle and turns it left to open. Handle Press: The agent grasps the handle and presses it downwards.
These tasks cover a range of manipulation skills, including pushing, pulling, grasping, placing, pressing, and turning, and they involve different objects and goals, providing a comprehensive evaluation of the agent’s ability to learn and generalize across various manipulation tasks.
Figure 5: Examples of 11 tasks in meta-world
Appendix D Dataset Construction
We construct our dataset from expert demonstrations collected in the Meta-World (Yu et al. 2020) simulator. For each task, successful execution trajectories are recorded and stored as NPZ and PKL files. Each trajectory contains RGB, depth, and mask images at each timestep, along with task metadata such as environment name, frame index, and camera parameters.
To enhance the usefulness of the dataset for learning fine-grained control, we identify fine manipulation intervals using a spatial-aware strategy. Specifically, we compute the 3D positions of both grippers and task-relevant target points by combining mask and rgb-depth information into point clouds. Distances between grippers and targets are monitored across frames to identify close-contact phases indicative of fine manipulation.
[…]
  • Interval: sampling interval between frames
  • distance_threshold: maximum distance to consider a frame as fine-grained
  • consecutive frames: number of consecutive frames required to enter fine-grained phase
  • recovery needed frames: number of coarse frames required before re-entering fine phase
  • suppress single spike: whether to ignore short spikes of distance violations
  • max allowed anomaly: allowed anomaly frames within fine-grained segment
To improve spatial focus, we crop each RGB frame based on task-specific visual layouts. Depending on the task and viewpoint, center or specific regions are retained using manually determined crop windows. The final dataset consists of (1) sampled RGB/mask/depth frames, (2) frame-wise 2D/3D positions of grippers and targets, (3) their pairwise distances, and (4) spatial plan table in textual form.
Appendix E VLM Prompt Design
Figure 6: Detailed Design of ours VLM Prompt
We explore the use of vision-language models (VLMs) to provide structured feedback in both generation validation and action refinement. While prior works often use open-ended prompts to guide action planning (e.g., “How should the robot complete this task?”), we find that such free-form task descriptions or scene summaries yield inconsistent or inaccurate responses in our setting.
Ineffective Prompt Types (Abandoned):
  • Task description (ineffective): Prompts like “close the door slow” or “grasp the basketball carefully” often fail to produce actionable or reliable feedback due to ambiguity and lack of spatial grounding.
  • Scene summary (ineffective): Descriptions of the visual scene, such as “the door is slightly open, and the gripper is nearby”, result in inconsistent VLM interpretation and lack of precise spatial understanding.
Effective Prompt Design:
We instead adopt a structured prompt format tailored to two specific functionalities shown in Fig. 6:
  1. 1.
    Video validation (generation filtering): We use a closed-loop prompt to determine whether a generated video should be accepted or rejected. The prompt emphasizes object completeness and task progression, while tolerating minor visual artifacts. The VLM is instructed to respond with a strict “Accept” or “Reject” based on visual quality.
    This strict, rule-based instruction significantly improves agreement with human judgments, especially in detecting object disappearance or hallucination.
  2. 2.
    Spatial plan table generation: As described in Section  D, we format spatial plan table into structured multi-step plans of the form:
    1Plan:
    21. move [-1, 0, 0] [0.19]
    32. move [0, 0, -1] [0.21]
    43. push [0, 0, 0] [0.00]
    Each line encodes an action, a direction vector (x: right-/left+, y: up+/down-, z: forward-/back+), and a positive float distance (in meters). These are generated by aligning frame transitions with gripper/target spatial dynamics and then converted into text for conditioning our VLM-enhanced modules.
Conclusion:
Our experiments suggest that free-form prompts are unreliable for task-level planning or video validation in closed-loop robotic control. Instead, spatial plan table with explicit roles, response formats, and spatial semantics lead to significantly more stable and interpretable VLM outputs, which are essential for real-time feedback in visual policy generation.
Appendix F Frame Matching Mechanism
To ensure precise alignment between visual goals and execution states, we introduce a similarity-based frame matching mechanism. At each timestep, the current observation frame $\mathbf{I}_{\text{cur}}$ is compared to the current target frame $\hat{\mathbf{I}}_{g}$ sampled from the predicted video sequence $\{\hat{\mathbf{I}}_{1},\hat{\mathbf{I}}_{2},\ldots,\hat{\mathbf{I}}_{K}\}$. We compute a composite similarity score via:
$\begin{split}\text{Sim}(\mathbf{I}_{\text{cur}},\hat{\mathbf{I}}_{g})=&\ w_{\text{geo}}\cdot\text{Sim}_{\text{geo}}+w_{\text{pos}}\cdot\text{Sim}_{\text{pos}}\\ &+w_{\text{ssim}}\cdot\text{SSIM}(\mathbf{I}_{\text{cur}},\hat{\mathbf{I}}_{g})+w_{\text{flow}}\cdot\text{Sim}_{\text{flow}},\end{split}$ (7)
where $\text{Sim}_{\text{geo}}$ is the similarity based on contour and shape features (Lowe 2004), $\text{Sim}_{\text{pos}}$ encodes block-wise intensity statistics and center-of-mass shifts, SSIM is the structural similarity index between grayscale frames (Wang et al. 2004), and $\text{Sim}_{\text{flow}}$ measures the inverse magnitude of optical flow (Farnebäck 2003). The weights are set to $w_{\text{geo}}{=}0.35$, $w_{\text{pos}}{=}0.35$, $w_{\text{ssim}}{=}0.2$, and $w_{\text{flow}}{=}0.1$. A frame is considered matched if:
$\text{Sim}(\mathbf{I}_{\text{cur}},\hat{\mathbf{I}}_{g})\geq\tau,$ (8)
with the threshold $\tau{=}0.8$. Upon match, we advance the goal frame to $\hat{\mathbf{I}}_{g+1}$. To prevent deadlocks when similarity fails to exceed the threshold, a hard switch is enforced every $T_{\text{max}}{=}28$ steps.
Appendix G Robustness to Occlusions via Masking
Figure 7: Video generation results with partial occlusion by random masking. Despite masked regions (left), the model generates coherent future frames (right) and maintains correct task execution.
To evaluate the robustness of our video generation model under partial visual occlusions, we introduce a masking mechanism before feeding the current frame into the diffusion model. Specifically, we randomly mask a portion of the image input, simulating occlusions such as those caused by environmental factors or sensor noise.
As illustrated in Figure 7, although significant areas of the current frame are masked, the generated frames remain visually consistent and the downstream agent successfully completes the manipulation tasks. This indicates that the model has learned a robust representation of both visual and temporal context, and can infer plausible continuations even when critical input regions are missing.

### v2

@@
 Abstract
-Vision-centric hierarchical embodied models have demonstrated strong potential for long-horizon robotic control.
+Vision-centric hierarchical embodied models have demonstrated strong potential.
 However, existing methods lack spatial awareness capabilities, limiting their effectiveness in bridging visual plans to actionable control in complex environments.
 To address this problem, we propose Spatial Policy (SP), a unified spatial-aware visuomotor robotic manipulation framework via explicit spatial modeling and reasoning.
-Specifically, we first design a spatial-conditioned embodied video generation module to model spatially guided predictions through a spatial plan table.
-Then, we propose a spatial-based action prediction module to infer executable actions with coordination.
+Specifically, we first design a spatial-conditioned embodied video generation module to model spatially guided predictions through the spatial plan table.
+Then, we propose a flow-based action prediction module to infer executable actions with coordination.
 Finally, we propose a spatial reasoning feedback policy to refine the spatial plan table via dual-stage replanning.
-Extensive experiments show that SP significantly outperforms state-of-the-art baselines, achieving a 33.0% average improvement over the best baseline.
-With an 86.7% average success rate across 11 diverse tasks, SP substantially enhances the practicality of embodied models for robotic control applications.
+Extensive experiments show that SP substantially outperforms state-of-the-art baselines, achieving over 33% improvement on Meta-World and over 25% improvement on iTHOR, demonstrating strong effectiveness across 23 embodied control tasks.
+We additionally evaluate SP in real-world robotic experiments to verify its practical viability.
+SP enhances the practicality of embodied models for robotic control applications.
 Code and checkpoints are maintained at https://plantpotatoonmoon.github.io/SpatialPolicy/.
-Introduction
-In recent years, hierarchical embodied models combining high-level planning and low-level action have achieved remarkable success in long-horizon robotic manipulation (Brohan et al.
-2024; Cheang et al.
-2024; Ze et al.
-2024; Li et al.
-2025b).
+1 Introduction
+In recent years, hierarchical embodied models combining high-level planning and low-level action have achieved remarkable success in robotic manipulation [rt2, gr2, dp3d, spvla].
 Among these approaches, vision-based methods, known as visuomotor robotic manipulation, have emerged as a particularly promising direction.
-By leveraging rich visual cues, these methods enable effective cross-hierarchical information transfer across diverse robotic embodiments and environments (Du et al.
-2023; Tian et al.
-2024; Ko et al.
-2023; Soni et al.
-2025; Black et al.
-2023; Li et al.
-2025a).
-Specicially, visuomotor robotic manipulation employs a generative model to predict trajectories of future video, which are subsequently translated into executable actions through a action prediction module.
-Despite significant advances, existing visuomotor robotic manipulation frameworks lack spatial awareness ability, failing to effectively bridge visual plans to executable actions in complex environments (Tian et al.
-2024; Black et al.
-2023).
-For instance, AVDC (Ko et al.
-2023) lacks explicit spatial modeling in its video generation process, resulting in physically implausible predictions, such as robotic arms passing through walls.
-Feedback module of CLOVER (Bu et al.
-2024) enhances dynamic environment adaptation, but its inability to process spatial relationships prevents handling of occluded target scenarios (Tian et al.
-2024).
+By leveraging rich visual cues, these methods enable effective cross-hierarchical information transfer across diverse robotic embodiments and environments [unipi, seer, avdc, videoagent, susie, grmg].
+Specifically, visuomotor robotic manipulation employs a generative model to predict trajectories of future video, which are subsequently translated into executable actions through an action prediction module.
+Despite significant advances, existing visuomotor robotic manipulation frameworks lack spatial awareness abilities, failing to effectively bridge visual plans to executable actions in complex environments [seer, susie].
+For instance, AVDC [avdc] lacks explicit spatial modeling in its video generation process, resulting in physically implausible predictions, such as robotic arms passing through walls.
+Feedback module of CLOVER [clover] enhances dynamic environment adaptation, but its inability to process spatial relationships prevents handling of occluded target scenarios [seer].
 To tackle this problem, we investigate the spatial-aware visuomotor robotic manipulation via explicit spatial modeling and reasoning, which introduces three key challenges: (1) how to generate videos that explicitly model spatial relations, such as object-goal positioning and task alignment; (2) how to utilize the generated videos to guide action prediction in a way that preserves spatial consistency and task relevance; and (3) how to reason over spatial cues during execution to enable real-time feedback that effectively bridges visual and control.
@@
 The core idea is to unify spatial modeling and reasoning via a spatial plan table that guides video generation, enables spatially consistent action prediction, and supports feedback-driven refinement during execution.
-Specifically, (1) we propose a Spatial-Conditioned Embodied Video Generation module that generates visual predictions guided by a structured spatial plan table, which encodes object-goal relations and directional cues to ensure spatial coherence in imagined futures, addressing the challenge of modeling spatially meaningful video trajectories; (2) we introduce a Spatial-Based Action Prediction module that infers executable actions by capturing spatial motion dynamics between predicted frames using spatial coordinates, tackling the challenge of translating visual predictions into spatially consistent actions; (3) we develop a Spatial Reasoning Feedback Policy that refines the spatial plan and video execution through dual-stage replanning based on VLM feedback and rule-based corrections, solving the challenge of real-time spatial reasoning during policy execution.
-Extensive experiments demonstrate that our model outperforms existing baselines, achieving 86.7% performance.
+Specifically, (1) we propose a Spatial-Conditioned Embodied Video Generation module that generates visual predictions guided by a structured spatial plan table, which encodes object-goal relations and directional cues to ensure spatial coherence in imagined futures, addressing the challenge of modeling spatially meaningful video trajectories; (2) we introduce a Flow-Based Action Prediction module that infers executable actions by capturing spatial motion dynamics between predicted frames using flow and spatial coordinates, tackling the challenge of translating visual predictions into spatially consistent actions; (3) we develop a Spatial Reasoning Feedback Policy that refines the spatial plan and video execution through dual-stage replanning based on VLM feedback and rule-based corrections, solving the challenge of real-time spatial reasoning during policy execution.
+Extensive experiments show that our model surpasses existing baselines, achieving 86.7% performance on Meta-World and 59.6% on iTHOR and further proving its viability through real-world robotic experiments.
 The contributions of this paper are summarized as follows:
@@
   • Modular design with structured spatial modeling and reasoning.
-SP consists of three synergistic modules: a spatial-conditioned video generation module guided by a structured spatial plan table, a spatial-based action prediction that preserves spatial consistency, and a spatial reasoning feedback policy that refines execution through real-time spatial replanning.
+SP consists of three synergistic modules: a spatial-conditioned video generation module guided by a structured spatial plan table, a flow-based action prediction that preserves spatial consistency, and a spatial reasoning feedback policy that refines execution through real-time spatial replanning.
   • Comprehensive evaluation and strong performance.
 We conduct extensive experiments on the MetaWorld benchmark to assess the impact of spatial integration, including ablations on video generation, action prediction, and feedback mechanisms.
-SP outperforms prior methods by over 30%, demonstrating the effectiveness of spatial reasoning in visuomotor robotic manipulation.
-Related Work
-Video Generation for Embodied Manipulation
+SP yields an absolute improvement of over 33% on Meta-World and over 25% on iTHOR and futher validate our model on real-world robotic executions, demonstrating the effectiveness of spatial reasoning in visuomotor robotic manipulation.
+2 Related Work
+2.1 Video Generation for Embodied Manipulation
 Video generation has recently been explored as a visual planning mechanism for embodied agents.
-AVDC (Ko et al.
-2023) generates action-annotated video plans from raw RGB inputs using dense correspondencesenabling policy learning without action supervision.
+AVDC [avdc] generates action-annotated video plans from raw RGB inputs using dense correspondencesenabling policy learning without action supervision.
 However, due to the lack of explicit spatial reasoning in both generation and action prediction, limiting precision in manipulation tasks.
-VideoAgent (Soni et al.
-2025) refines video plans via self-conditioning to reduce hallucinations and improve visual quality.
+VideoAgent [videoagent] refines video plans via self-conditioning to reduce hallucinations and improve visual quality.
 However, it inherits the lack of explicit spatial reasoning from AVDC in both generation and action prediction, limiting precision in geometry-sensitive manipulation tasks.
-Clover (Bu et al.
-2024) introduces a closed-loop framework combining video diffusion planning and feedback control via an embedding-based error signal, achieving strong results in long-horizon tasks.
+Clover [clover] introduces a closed-loop framework combining video diffusion planning and feedback control via an embedding-based error signal, achieving strong results in long-horizon tasks.
 However, its spatial modeling is limited to depth-based cues and lacks explicit 3D spatial planning or structured geometric reasoning, which restricts generalization across diverse spatial layout.
-SEER (Tian et al.
-2024)proposes a Predictive Inverse Dynamics Model (PIDM) that forecasts future visual states and infers actions end-to-end, closing the loop between vision and control.
+SEER [seer] proposes a Predictive Inverse Dynamics Model (PIDM) that forecasts future visual states and infers actions end-to-end, closing the loop between vision and control.
 While it achieves strong performance through large-scale training, it lacks explicit spatial reasoning modules and does not leverage structured spatial plans, which may limit interpretability and adaptability in geometry-sensitive tasks.
-SuSIE (Black et al.
-2023) leverages an image-editing diffusion model to generate visual subgoals for a low-level goal-conditioned policy, enabling better generalization in unstructured environments.
+SuSIE [susie] leverages an image-editing diffusion model to generate visual subgoals for a low-level goal-conditioned policy, enabling better generalization in unstructured environments.
 However, its subgoal generation relies on appearance-based edits without explicit incorporation of 3D spatial relationships, limiting its precision in spatially complex manipulation scenarios.
-Spatial Reasoning in Vision-Language Embodied Systems
+2.2 Spatial Reasoning in Vision-Language Embodied Models
 Spatial understanding has been addressed in embodied AI from several perspectives.
-SpatialVLM (Chen et al.
-2024a) and GraspCorrect (Lee, Hong, and Kim 2025) leverage pretrained vision-language models to extract chain-of-thought spatial cues and determine object relations in manipulation tasks.
-3D-VLA (Zhen et al.
-2024) explicitly encodes spatial semantics by aligning 3D vision with textual instructions.
-SPA (Zhu et al.
-2025) utilizes point cloud reconstruction for spatial alignment in manipulation and navigation tasks.
-Meanwhile, LLM-Planner (Song et al.
-2023)and related methods employ large language models to infer spatially structured plans in symbolic form.
-Despite these advances, such methods either focus on semantic spatial understanding or operate in symbolic domains, lacking visual feedback and executable rollout.
-Moreover, most rely on handcrafted prompts or static spatial priors and do not support continuous video generation.
-In contrast, our work closes this gap by embedding structured spatial plans—automatically derived from environment geometry—into a generative video prediction framework, thereby tightly coupling spatial reasoning and visual synthesis for action-aware planning.
-Spatial Conditioning in Generative Video Models
-Recent advances in text-to-video diffusion models, such as AnimateDiff (Guo et al.
-2024) and VideoCrafter2 (Chen et al.
-2024b), achieve high-fidelity generation but lack controllability over agent behavior or scene dynamics.
-ScenarioDiff (Zhang et al.
-2025) and DirecT2V (Hong et al.
-2024) propose spatially guided generation using manually specified constraints like bounding boxes or motion paths.
-These constraints are often synthetic and decoupled from real embodied agent states, limiting their adaptability and semantic alignment with embodied control tasks.
-Spatial Policy
+SpatialVLM [spatialvlm] and GraspCorrect [graspcorrect] leverage pretrained vision-language models to extract chain-of-thought spatial cues and determine object relations in manipulation tasks.
+3D-VLA [3dvla] explicitly encodes spatial semantics by aligning 3D vision with textual instructions.
+SPA [spa] utilizes point cloud reconstruction for spatial alignment in manipulation and navigation tasks.
+Meanwhile, LLM-Planner [llmplanner] and related methods employ large language models to infer spatially structured plans in symbolic form.
+Despite these advances, existing methods primarily address semantic or symbolic spatial reasoning and lack visual feedback coupled with executable rollouts.
+Many rely on handcrafted prompts or fixed spatial priors and do not support continuous video prediction.
+Our approach fills this gap by embedding structured spatial plans, automatically derived from the environment geometry, into a generative video prediction framework, thereby integrating spatial reasoning and visual synthesis for action-aware planning.
+Figure 1: Framework Overview of Spatial Policy.
+Our system comprises three modules: (1) Spatial-Conditioned Embodied Video Generation, which uses a Spatial Plan Table obtained via VLM reasoning over the spatial offset between the robot and the object, to guide diffusion model in generating spatially coherent video prediction.
+(2) Flow-Based Action Prediction, which converts the generated video into executable actions using flow and spatial coordinates.
+(3) Spatial Reasoning Feedback Policy, which enables real-time correction via a dual-stage replanning strategy, combining VLM-based video judgement and policy diagnostics to refine the spatial plan table for closed-loop control.
+3 Spatial Policy
 We propose a visuomotor robotic manipulation framework that integrates structured spatial reasoning into the perception-to-control pipeline, enabling robots to generate spatially consistent visual plans and execute them reliably.
@@
 1, our system comprises three interconnected modules.
-First, the Spatial-Conditioned Embodied Video Generation module generates future video trajectories conditioned on a structured spatial plan table composed of atomic actions, directional vectors, and relative distances.
-Second, the Spatial-Based Action Prediction module predicts actions aligned with the generated video using the spatial coordinates.
+First, the Spatial-Conditioned Embodied Video Generation module in Sec.
+3.1 generates future video trajectories conditioned on a structured spatial plan table composed of atomic actions, directional vectors, and relative distances.
+Second, the Flow-Based Action Prediction module in Sec.
+3.2 predicts actions aligned with the generated video using the flow and spatial coordinates.
 Third, the Spatial Reasoning Feedback Policy monitors spatial alignment during execution and dynamically refines both the spatial plan table and the visual predictions.
 Together, these modules form a closed-loop framework that maintains spatial consistency across video imagination, action prediction, and feedback-based reasoning.
-Figure 1: Framework Overview of Spatial Policy.
-Our system comprises three modules: (1) Spatial-Conditioned Embodied Video Generation, which uses a Spatial Plan Table obtained via VLM reasoning over the spatial offset between the robot and the object, to guide diffusion model in generating spatially coherent video prediction.
-(2) Spatial-Based Action Prediction, which converts the generated video into executable actions using spatial coordinates.
-(3) Spatial Reasoning Feedback Policy, which enables real-time correction via a dual-stage replanning strategy, combining VLM-based video judgement and policy diagnostics to refine the spatial plan table for closed-loop control.
-Spatial Conditioned Embodied Video Generation
+The full inference pipeline is provided in Appendix A.1.
+3.1 Spatial-Conditioned Embodied Video Generation
 To generate spatially grounded visual trajectories for visuomotor robotic manipulation, we adopt a conditional diffusion framework for future video prediction.
 The model takes as input an initial observation frame and a textual task description, and generates a sequence of future frames representing high-level visual plans.
-To enhance spatial awareness, we additionally introduce spatial plan table derived from spatial relationships as an auxiliary global condition.
+To enhance spatial awareness, we additionally introduce the spatial plan table derived from spatial relationships as an auxiliary global condition.
 Spatial Plan Table
@@
 1.
+The detailed VLM inference procedure is provided in Appendix A.5.
 Encoding and Integration
@@
 $\mathbf{e}_{\text{subplan}}=\text{MLP}\left([\mathbf{e}_{\text{act}};\,\mathbf{e}_{\text{dir}};\,\mathbf{e}_{\text{dis}}]\right),$ (1)
-All subgoal embeddings $\{\mathbf{e}_{1},\dots,\mathbf{e}_{n}\}$ are concatenated with the task text embedding $\mathbf{z}_{\text{task}}\in\mathbb{R}^{d}$ obtained via CLIP (Radford et al.
-2021) to form the composite global condition:
+All subgoal embeddings $\{\mathbf{e}_{1},\dots,\mathbf{e}_{n}\}$ are concatenated with the task text embedding $\mathbf{z}_{\text{task}}\in\mathbb{R}^{d}$ obtained via CLIP [clip] to form the composite global condition:
 $\mathcal{T}=\left[\mathbf{z}_{\text{task}};\,\mathbf{e}_{1};\,\dots;\,\mathbf{e}_{n}\right].$ (2)
@@
 Conditional Diffusion Formulation
-We follow the design of a u-net-based denoising diffusion probabilistic model, adapted from the Imagen architecture (Ho et al.
-2022), consistent with AVDC (Ko et al.
-2023).
+We follow the design of a U-Net-based denoising diffusion probabilistic model, adapted from the Imagen architecture [imagen], consistent with AVDC [avdc].
 Given an initial frame $\mathbf{I}_{0}\in\mathbb{R}^{H\times W\times 3}$ and a composite condition vector $\mathcal{T}$ that incorporates both task instructions and structured spatial information, the model predicts a sequence of $T=7$ future frames $\mathbf{I}_{1:T}$, resulting in an 8-frame trajectory.
@@
 where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is sampled noise, $\beta_{t}$ is the diffusion schedule, and $\epsilon_{\theta}$ is the model’s denoising network.
-Further training details are provided in the Appendix.
-Spatial Based Action Prediction
-To convert spatially grounded visual plans into executable actions, we design an Spatial Coordinate-Guided Action Prediction module built upon the diffusion policy framework (Chi et al.
-2023; Ji et al.
-2025), due to its strong capacity for modeling multimodal distributions and complex temporal dependencies, making it particularly well-suited for capturing the spatial variability and uncertainty inherent in visual planning.
-The model predicts multi-step actions that guide the robot from the current observation toward the visual goal implicitly defined by the generated video sequence, leveraging spatial coordinate to capture fine-grained spatiotemporal dynamics.
+3.2 Flow-Based Action Prediction
+To incorporate explicit motion cues into action generation, we extend the original module by adding flow as an additional conditioning signal, while preserving the overall framework and notation style.
 Multimodal Spatial Input Encoding
-We condition our policy on three types of information: (1) the current image $\mathbf{I}_{\text{cur}}$, (2) the goal image $\mathbf{I}_{\text{goal}}$ (sampled from the generated visual plan within the next $k$ steps, i.e., $\mathbf{I}_{\text{goal}}=\mathbf{I}_{t+\delta}$ where $1\leq\delta\leq k$), and (3) the spatial coordinate of robotic arms $\mathbf{p}_{\text{cur}}\in\mathbb{R}^{3}$.
-These inputs capture both spatial and temporal motion cues.
-All inputs are encoded into compact feature vectors: the image pair is stacked along the channel dimension and passed through a shared ResNet encoder, while the spatial coordinate is encoded via a shallow MLP.
-The resulting modality-specific features are then concatenated to form a unified condition vector, which guides the denoising network in our diffusion policy-based action prediction model.
-Diffusion Policy-based Action Prediction Model
-We formulate action prediction as a conditional generative process. Specifically, we adopt a unet-based diffusion policy architecture to model the conditional distribution over action trajectories:
-$p_{\theta}(\mathbf{a}_{0:H-1}\mid\mathbf{I}_{\text{cur}},\mathbf{I}_{\text{goal}},\mathbf{p}_{\text{cur}})$ (5)
+We condition our policy on four types of information: (1) the current image $\mathbf{I}_{\text{cur}}$, (2) the goal image $\mathbf{I}_{\text{goal}}$ (sampled from the generated visual plan within the next $k$ steps), (3) the spatial coordinate of the robotic arm $\mathbf{p}_{\text{cur}}\in\mathbb{R}^{3}$, and (4) the dense flow field computed between consecutive frames.
+The flow field is computed using the classical Farnebäck method [farneback]:
+$\mathbf{F}_{t}=\text{Flow}\!\left(\mathbf{I}_{t-1},\mathbf{I}_{t}\right),$ (5)
+and then encoded using a CNN-based encoder.
+All inputs are encoded into compact feature vectors through their respective encoders (ResNet for images and an MLP for spatial coordinates).
+The modality-specific features are subsequently concatenated to form the global condition vector used by the diffusion policy.
+Diffusion Policy-based Action Prediction Module
+We formulate action prediction as a conditional generative process. Specifically, we adopt a U-net-based diffusion policy architecture to model the conditional distribution over action trajectories:
+$p_{\theta}(\mathbf{a}_{0:H-1}\mid\mathbf{I}_{\text{cur}},\mathbf{I}_{\text{goal}},\mathbf{p}_{\text{cur}},\mathbf{F}_{t})$ (6)
 where $\mathbf{a}_{0:H-1}$ denotes a sequence of $H$ actions, each including a 3-DoF end-effector pose and gripper control. We train the policy via denoising score matching, minimizing:
 $\mathcal{L}_{\text{MSE}}=\mathbb{E}_{\mathbf{a},\epsilon,t}\left\|\epsilon-\epsilon_{\theta}\!\left(\sqrt{1-\beta_{t}}\cdot\mathbf{a}+\sqrt{\beta_{t}}\cdot\epsilon,t\mid\mathbf{c}\right)\right\|^{2}$ (7)
-where $\mathbf{c}$ denotes the concatenated global condition vector from all encoded inputs, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is Gaussian noise, and $t$ is the diffusion timestep.
+where $\mathbf{c}$ denotes the concatenated global condition vector including all encoded modalities, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is Gaussian noise, and $t$ is the diffusion timestep.
 $\epsilon_{\theta}$ is the learned denoising network.
 At inference time, the model predicts the entire action trajectory in a denoising pass, and the resulting sequence is executed step-by-step in the environment.
-Further training details are provided in the Appendix.
-Spatial Reasoning Feedback Policy
+The frame–switching mechanism used during inference is described in Appendix A.4.
+3.3 Spatial Reasoning Feedback Policy
 To ensure spatial consistency throughout execution, we introduce a Spatial Reasoning Feedback Policy that enables closed-loop correction by dynamically refining the Spatial Plan Table based on perceptual feedback.
 This module integrates a dual-stage feedback mechanism, operating at both the video validation and action execution levels.
-During visual planning, we incorporate a vision-language model (VLM)-based consistency checker inspired by VideoAgent (Soni et al.
-2025).
+During visual planning, we incorporate a vision-language model (VLM)-based consistency checker inspired by VideoAgent [videoagent].
 A pretrained GPT-4o model is prompted to assess the spatial plausibility of the generated video—verifying object-goal alignment, manipulator trajectories, and overall spatial coherence.
@@
 Through this real-time refinement of spatial intent and control guidance, our system maintains alignment between visual imagination and robotic execution, enabling robust long-horizon manipulation.
-Experiments
-Method door-open door-close basketball shelf-place btn-press btn-press-top
-AVDC 30.7% 28.0% 21.3% 8.0% 34.7% 17.3%
-AVDC-R 72.0% 89.3% 37.3% 18.7% 60.0% 24.0%
-VA 40.0% 29.3% 13.3% 9.3% 38.7% 18.7%
-VA-OR 82.7% 97.3% 40.0% 26.7% 73.3% 44.0%
-SP 98.6% 100.0% 66.7% 80.0% 94.7% 96.0%
-SP-R 98.6% 100.0% 73.3% 92.0% 96.0% 98.6%
-Method faucet-close faucet-open handle-press hammer assembly Overall
-AVDC 12.0% 17.3% 41.3% 0.0% 5.3% 19.6%
-AVDC-R 53.3% 24.0% 81.3% 8.0% 6.7% 43.1%
-VA 46.7% 12.0% 36.0% 0.0% 1.3% 22.3%
-VA-OR 74.7% 46.7% 86.7% 8.0% 10.7% 53.7%
-SP 93.3% 89.3% 100.0% 74.7% 24.0% 83.4%
-SP-R 92.0% 96.0% 100.0% 73.3% 32.0% 86.7%
+4 Experiments
 Table 1: Comparison of task success rates (%) across 11 MetaWorld tasks.
@@
 Bold indicates best.
-Datasets and Experimental Setups
+Method door-open door-close basketball shelf-place btn-press btn-press-top
+AVDC [avdc] 30.7% 28.0% 21.3% 8.0% 34.7% 17.3%
+AVDC-R [avdc] 72.0% 89.3% 37.3% 18.7% 60.0% 24.0%
+VA [videoagent] 40.0% 29.3% 13.3% 9.3% 38.7% 18.7%
+VA-OR [videoagent] 82.7% 97.3% 40.0% 26.7% 73.3% 44.0%
+SP 100.0% 100.0% 77.3% 78.7% 96% 98.6%
+SP-R 98.6% 100.0% 73.3% 92.0% 96.0% 98.6%
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 2: Comparison of task success rates (%) across 12 tasks in four rooms on iTHOR.
+Room AVDC [avdc] VideoAgent [videoagent] SP (Ours)
+Kitchen 26.7% 28.3% 88.3%
+Living Room 23.3% 26.7% 50.0%
+Bedroom 38.3% 41.7% 63.3%
+Bathroom 36.7% 40.0% 36.7%
+Average 31.3% 34.2% 59.6%
+In this section, we comprehensively evaluate our method across both manipulation and navigation tasks.
+Sec.
+4.1 introduces the environments and datasets, and Sec.
+4.2 reports the main results on Meta-World and iTHOR.
+Sec.
+4.3 presents the real-world evaluation results.
+We further conduct five ablation studies in Sec.
+4.4, examine failure modes in Sec.
+4.5, analyze computational efficiency in Sec.
+4.2, assess robustness to occlusions in Sec.
+4.6, and study key hyperparameters in Sec.
+4.7.
+4.1 Datasets and Experimental Setups
 Datasets and Environments.
-We follow the evaluation setup of (Ko et al.
-2023) and conduct experiments in the Meta-World (Yu et al.
-2020).
-Meta-World is a simulated benchmark designed for assessing robotic manipulation capabilities, consisting of 11 representative single-arm tasks (such as pushing, pulling, grasping, and inserting) performed by a Sawyer robotic arm.
-We run 25 random seeds across 11 tasks and 3 camera views, yielding over 800 rollouts in total, and evaluate 5 ablation studies in our model.
+We follow the evaluation setup of [avdc] and conduct experiments in both Meta-World [metaworld] and iTHOR [kolve2017ai2thor].
+Meta-World is a simulated robotic manipulation benchmark consisting of 11 single-arm control tasks (e.g., pushing, pulling, grasping, inserting); we run 25 random seeds across 11 tasks and 3 camera views, yielding over 800 rollouts.
+iTHOR is an interactive 3D household environment featuring four room types and 12 navigation tasks, where each task is evaluated with 20 random seeds to assess robustness and cross-scene generalization.
+Complete training details are provided in Appendix A.2, generated results in Appendix A.3, and dataset details in Appendix A.4.
 Baselines and Our Methods
-  • AVDC: A two-stage control planning framework is employed, wherein a pretrained diffusion-based generative model synthesizes a video sequence, followed by the estimation of inter-frame motion using an optical flow network.
-The resulting motion is then used to inversely infer the corresponding action sequence.
-  • AVDC-Replan: An extension of AVDC that incorporates a replanning mechanism.
-During execution, if action failure or robotic arm stagnation is detected, a new video is regenerated and a revised action plan is inferred.
-  • VideoAgent: This method employs a self-supervised consistency mechanism to iteratively refine the generated video.
-Multiple candidate video plans are scored using GPT-4, and the most promising one is selected for execution.
-  • VideoAgent-online-Replan: An extension of VideoAgent that continuously collects successful execution trajectories during deployment.
-These are used to update both the video generation and optimization models online.
-A replanning strategy is also introduced to regenerate videos when execution failures occur.
-  • SP (Ours): We extract structured planning information from spatial context using VLM, and inject it into the diffusion model via global condition to effectively guide video generation.
-The generated video is subsequently translated into executable actions using a goal-conditioned policy.
-  • SP-Replan (Ours): An extended version of our method that performs iterative video generation in the planning phase.
-During execution, if robotic arm stagnation is detected, a new video is regenerated and the action sequence is re-inferred.
-Main Results
-As shown in Table 1, our method achieves a success rate of 83.4% without requiring execution-time replanning, already outperforming all prior baselines.
-In contrast, methods such as AVDC and VA rely on task text to generate videos or pretrained optical flow to compute displacement coordinates for action prediction.
-These baselines are fragile during deployment, once a deviation or unexpected event occurs, they must fall back on explicit replanning to recover; otherwise, the robot stalls.
-By contrast, our approach leverages spatial plans to generate structured, goal-aligned videos and predicts executable actions that inherently adapt to spatial variations, leading to more robust and consistent task completion.
-SP-R further boosts performance to 86.7%, demonstrating the effectiveness of incorporating real-time spatial reasoning into the control loop.
-On the subset of challenging tasks ( Assembly, Hammer, Faucet Open, Button Press Top, Shelf Place, and Basketball ), our method achieves an average success rate of 77.5%, significantly outperforming the best-performing baseline, which only reaches 29.2% .
-This corresponds to a relative improvement of 165.4%.
-On simpler tasks, our approach achieves near-perfect performance, often approaching 100% success.
-Figure 2: Comparative visualization of spatial planning and video generation results for two representative tasks: Shelf Place and Basketball.
-The yellow line depicts the motion trajectory; the symbol marks the initial position, the symbol indicates the final task position, and the symbols denote intermediate steps along the path.
-In the Shelf Place task, as shown in Figure 2, the baseline method generates an incorrect and inflexible trajectory (highlighted in the yellow box), resulting in task failure due to its inability to adapt to environmental variations.
-In contrast, our approach, enhanced with spatial understanding, enables the robotic arm to follow a dynamically planned path and successfully complete the placement task.
-Similarly, in the Basketball task (also illustrated in Figure 2), our method exhibits refined spatial reasoning by allocating more frames (highlighted in the green box) around the critical manipulation event—grasping the ball—leading to smoother and more precise motions.
-The baseline method, lacking this focused spatial adaptation, often misses essential interaction moments.
-These results underscore the benefits of integrating spatial awareness with dynamic feedback for robust and effective long-horizon robotic planning.
-Figure 3: Visualization of the replanning-based spatial planning process for a manipulation task.
-Figure 3 further demonstrates how our method leverages spatial feedback and iterative replanning to recover from failures and achieve successful task execution.
-Unlike baseline planning, which can fail when the initial path is invalid, our approach dynamically adapts, resulting in more robust and reliable performance.
-Ablation Study
-Ablations on the Video Generation Model.
-We compare three strategies in the video generation module to evaluate their impact on video quality and downstream control performance.
-(1) The baseline uses AVDC’s video generation approach, where a diffusion model is conditioned only on task text and the initial frame.
-(2) The Distance variant directly encodes the spatial offset between the robot and target object (spatial coordinates) and concatenates it with the task prompt as input.
-(3) The Subplan variant leverages a vision-language model to transform spatial coordinates into structured subplans that encode action type, direction, and distance, which are then injected into the diffusion model.
-As shown in Table 2, both spatial variants outperform the AVDC baseline, demonstrating the utility of explicitly modeling spatial structures for enhanced visuomotor performance.
+  • AVDC: Generates videos via a pretrained diffusion model, extracts optical-flow motion, and inversely infers actions.
+  • AVDC-Replan: Adds replanning by regenerating videos when failures or stagnation occur.
+  • VideoAgent: Refines video plans with a self-supervised consistency rule and selects candidates via GPT-4 scoring.
+  • VideoAgent-online-Replan: Online-updates video generation and optimization using successful trajectories and triggers replanning upon failures.
+  • SP (Ours): Extracts structured spatial plans with a VLM and injects them as global conditions into the diffusion model; actions are derived via a goal-conditioned policy.
+  • SP-Replan (Ours): Adds iterative regeneration and replanning when execution stagnation is detected.
+4.2 Main Results
+Meta-World
+As shown in Table 1, our method achieves an overall success rate of 86.7%, outperforming all prior baselines.
+Approaches such as AVDC and VA rely on task text or pretrained flow to infer motion, making them brittle under execution drift and forcing explicit replanning when deviations occur.
+In contrast, our spatial-plan-driven framework produces structured, goal-aligned video predictions and actions that inherently adapt to spatial variations, resulting in more stable and consistent execution.
+On the subset of challenging tasks (Assembly, Hammer, Faucet Open, Button Press Top, Shelf Place, Basketball), our method reaches 77.5% success, whereas the best baseline achieves only 29.2%, corresponding to a 165.4% relative improvement.
+Easier tasks are solved nearly perfectly, often approaching 100% success.
+Figure 2: Comparisons on Shelf Place and Basketball in Meta-World. The yellow line shows the motion trajectory; marks the start, the final position, and intermediate steps.
+Figure 3: Comparisons on Painting and Mirror in iThor. The bounding box marks the target object.
+Fig.
+2 illustrates these advantages.
+In Shelf Place, the baseline produces an inflexible trajectory that cannot correct for scene variations, whereas our model generates adaptive, spatially coherent paths.
+In Basketball, our method allocates more frames to the critical grasping phase (green box), yielding smoother and more reliable manipulation, while baseline methods frequently miss key interaction moments.
+Figure 4: Example of iterative spatial replanning during manipulation in Meta-World.
+[…]
+iTHOR
+Table 2 shows that our method reaches an average success rate of 59.6% on iTHOR, clearly above the 31.3% of AVDC and the 34.2% of VideoAgent.
+The largest improvements appear in Kitchen with 88.3% and in Bedroom with 63.3%, where cluttered and variable layouts require precise reasoning about object positions and feasible motion.
+Living Room also benefits from stronger spatial grounding, reaching 50.0%.
+Bathroom performance is similar to the baselines due to its highly constrained geometry, which limits the impact of planning.
+Overall, the results indicate that explicit spatial structure, rather than relying solely on visual priors, is crucial for generalizing across diverse room configurations.
+The qualitative rollouts in Fig.
+3 further illustrate this advantage.
+Baseline trajectories drift because diffusion priors alone cannot preserve stable spatial alignment, leading to early pose errors in long-horizon tasks such as Painting and Mirror.
+This issue is amplified in iTHOR, where frequent camera rotations require the model to maintain scene consistency across changing viewpoints.
+Our method preserves the global layout more faithfully: generated frames remain coherent after rotation, and the target is localized consistently throughout the sequence.
+By encoding spatial plans that constrain depth reasoning, object–agent relations, and admissible motion, the model produces trajectories that better reflect the actual room topology even under viewpoint shifts.
+4.3 Real-World Experiments
+We evaluate our system in a real-world setting using an SO-100 robotic arm [so100_wiki] on a pick-doll task, where the robot must grasp a yellow rabbit toy and place it into a box.
+Since real-world data lack depth, we estimate depth with DepthAnything [depthanything] and feed the RGB frame, predicted depth, and object spatial information into the spatial plan table.
+For action prediction, we use the top-down current and goal frames together with the wrist-camera frame to improve spatial grounding.
+As shown in Figure 5, we evaluate 20 trials and achieve a 30% success rate, indicating that our model can still complete the task even under incomplete visual observations and depth estimation noise.
+Figure 5: Real-world execution of the pick doll task.
+4.4 Ablation Study
+Ablation on the Video Generation Module.
+We evaluate three strategies in the video generation module in Meta-World.
+(1) The Baseline follows AVDC and conditions the diffusion model only on the task prompt and initial frame.
+(2) Our Distance variant adds the spatial offset between the end effector and the target as raw spatial guidance.
+(3) Our Subplan variant transforms this offset into structured action cues using a vision language model and injects them as high level spatial guidance As shown in Table 3, both spatial variants outperform the AVDC baseline, demonstrating the utility of explicitly modeling spatial structures for enhanced visuomotor performance.
 Moreover, the subplan-based approach achieves further gains by abstracting raw spatial cues into structured guidance, highlighting the importance of structured spatial abstraction in facilitating precise and actionable visual plan generation.
-Ablations on the Action Prediction Model.
-AVDC uses a pre-trained flow model to estimate pixel-wise motion between the current and goal frames, then converts the flow into a spatial target coordinate.
-The robot executes actions using a classical controller to move toward the point and applies gripper control based on heuristics.
-In contrast, our method is built on diffusion policy and directly predicts executable actions from the current frame, goal frame, and robot spatial coordinates.
-This design captures spatial and temporal context jointly and allows flexible adaptation to unexpected events such as object shifts or failures.
-Results in Table 3 show that our approach achieves higher success rates and better robustness, confirming that learning-based spatial reasoning provides more effective action prediction than coordinate-based control.
+Ablation on the Action Prediction Module.
+We compare two action prediction strategies in Meta-World.
+(1) AVDC converts optical flow between the current and goal frames into a target coordinate and executes it with a classical controller and heuristic gripper rules.
+(2) Our method instead uses a flow-based diffusion policy that directly predicts actions from the current frame, goal frame, flow, and robot spatial coordinates, allowing spatial and temporal cues to be modeled jointly.
+As shown in Table 4, our flow-based diffusion policy consistently outperforms AVDC across all difficulty levels.
+The average success rate rises from 19.6% for AVDC to 63.4% for our method, with especially large gains on Medium, Hard, and Very Hard tasks.
+These results highlight that learning based spatial reasoning from flow yields more reliable and adaptable action generation than mapping flow fields to target coordinates through hand-designed controllers.
 Ablation on Spatial Condition Injection Strategy in Video Generation.
-To investigate how spatial conditions affect video generation, we conduct an ablation study comparing two injection strategies for integrating the spatial plan table into the diffusion model.
-(1) The spatial plan table is concatenated with the task embedding and propagated uniformly across all layers of the U-Net backbone in the Global Cond variant directly.
-This approach provides consistent spatial guidance throughout the generation process.
-(2) Inspired by GLIGEN (Li et al.
-2023), the spatial plan is injected only into intermediate U-Net layers via a cross-attention mechanism in the Local Cond variant, allowing localized spatial modulation while preserving the task-level prompt as a global condition.
-Both strategies are trained under identical configurations and evaluated on MetaWorld.
-As shown in Table 4, Local Conditioning slightly improves the average task success rate and shows better performance on harder tasks (e.g., +5.3% on Hard).
-These results suggest that injecting spatial structure via localized attention can enhance spatial precision during video synthesis, thereby improving policy success rates, especially in more challenging environments.
-Easy Medium Hard Very Hard Overall
+We compare two strategies for injecting spatial plans into the diffusion model.
+(1) The Global Cond variant concatenates the spatial plan table with the task embedding and applies it uniformly to all U-Net layers.
+(2)The Local Cond variant follows the GLIGEN formulation and injects the spatial plan only into intermediate layers through cross attention, enabling more localized spatial modulation.
+As shown in Table 5, Local Conditioning yields higher success rates, reaching 84.3% on average and improving performance on harder tasks such as Hard and Very Hard.
+This suggests that localized spatial injection improves spatial precision in video synthesis and leads to more reliable downstream control.
+Table 3: Comparison of task success rates (%) using different video generation strategies across four levels of task difficulty in Meta-World.
+AVDC uses text and initial frame only; Distance adds raw spatial offsets; Subplan injects structured spatial subplans.
+Easy Medium Hard Very Hard Average
 AVDC 75.0% 48.7% 26.7% 48.0% 63.4%
 Distance 93.5% 56.7% 25.3% 81.3% 79.4%
-Subplan 96.8% 63.3% 24.0% 78.7% 82.4%
-Table 2: Comparison of task success rates (%) using different video generation strategies across four levels of task difficulty.
-AVDC uses text and initial frame only; Distance adds raw spatial offsets; Subplan injects structured spatial subplans.
-Easy Medium Hard Very Hard Overall
-Flow 25.9% 10.7% 5.3% 8.0% 19.6%
+Subplan 94.8% 69.3% 21.3% 78.7% 82.0%
+Table 4: Comparison of task success rates (%) across four difficulty levels in Meta-World.
+AVDC refers to AVDC’s action predictor that directly maps optical flow to actions.
+Ours uses a flow-based diffusion policy.
+Easy Medium Hard Very Hard Average
+AVDC 25.9% 10.7% 5.3% 8.0% 19.6%
 Ours 75.0% 48.7% 26.7% 48.0% 63.4%
-Table 3: Comparison of task success rates (%) between our method and AVDC’s action prediction module across four levels of task difficulty.
-Flow denotes the flow model-based action prediction used in AVDC; Ours denotes our spatial-based action prediction module.
-Easy Medium Hard Very Hard Overall
-Global Cond 96.8% 63.3% 24.0% 78.7% 82.4%
+Table 5: Comparison of task success rates (%) using different spatial condition injection strategies for video generation across four levels of task difficulty in Meta-World.
+Global Cond injects the spatial plan table across all U-Net blocks; Local Cond follows GLIGEN-style partial injection via cross-attention at intermediate layers.
+Easy Medium Hard Very Hard Average
+Global Cond 94.8% 63.3% 21.3% 78.7% 82.0%
 Local Cond 99.0% 62.0% 29.3% 80.0% 84.3%
-Table 4: Comparison of task success rates (%) using different spatial condition injection strategies for video generation across four levels of task difficulty.
-Global Cond injects the spatial plan table across all U-Net blocks; Local Cond follows GLIGEN-style partial injection via cross-attention at intermediate layers.
-Ablations on the Replanning Mechanism.
-We compare three settings to evaluate the effect of the replanning mechanism: (1) no replanning at all, (2) replanning only during the video generation stage, and (3) replanning during both the video generation and real execution stages.
-As shown in Table 5, task success improves when replanning is introduced.
-Our experiments demonstrate that enabling replanning during both the imagination and real execution stages significantly enhances task success by facilitating dynamic adaptation to the evolving environment.
-While replanning solely during the imagination stage offers limited improvements, the integration of a vision-language-based evaluation module allows the system to selectively regenerate and retain higher-quality trajectories, effectively filtering out suboptimal predictions despite the inherent stochasticity of video generation.
-Without any replanning, the model’s performance deteriorates due to its inability to recover from unforeseen variations.
-Therefore, a combined replanning strategy is crucial for achieving robust and effective long-horizon robotic control.
-Easy Medium Hard Very Hard Overall
-NR 96.8% 63.3% 24.0% 78.7% 82.4%
+Ablation on the Replanning Mechanism.
+We compare three replanning settings: (1) no replanning, (2)replanning only during video generation, and (3) replanning during both video generation and real execution.
+As shown in Table 6, performance improves as more stages incorporate replanning.
+Replanning in both imagination and execution achieves the highest average success rate of 86.7%, with clear gains on harder tasks.
+While generation only replanning offers moderate improvement, combining it with execution time replanning enables the system to correct failures, adapt to environmental changes, and stabilize long horizon behavior.
+Without replanning, the policy cannot recover from unexpected variations.
+These results indicate that using replanning in both stages is essential for robust long horizon control.
+Table 6: Comparison of task success rates (%) using different replanning mechanisms across four levels of task difficulty in Meta-World.
+NR denotes no replanning; GR denotes generation replanning; ER denotes execution replanning.
+Bold indicates best.
+Easy Medium Hard Very Hard Average
+NR 94.8% 63.3% 21.3% 78.7% 82.0%
 GR 96.0% 70.7% 24.0% 80.0% 83.4%
 GR&ER 97.3% 73.3% 32.0% 92.0% 86.7%
-Table 5: Comparison of task success rates (%) using different replanning mechanisms across four levels of task difficulty.
-NR denotes no replanning; GR denotes generation replanning; ER denotes execution replanning.
-Bold indicates best.
-Ablations on Finetuning with Additional Execution Data.
-To assess the impact of trajectory diversity on policy robustness, we compare two training data configurations for our model.
-(1) The model is trained exclusively on demonstrations where tasks are completed successfully from start to finish without failure.
-(2) The model is further finetuned on an expanded dataset that includes recovery trajectories where execution initially failed but eventually succeeded when resumed from intermediate states.
-Finetuning on these recovery trajectories yields noticeable gains, particularly on challenging tasks.
-On btn-press-top, the success rate improves from 97.3% to 99.0%.
-More strikingly, on the more complex hammer task, performance increases from 68.0% to 78.7%.
-These results indicate that including partial-to-successful executions enhances the model’s robustness by exposing it to diverse recovery scenarios.
-Such data enables better generalization to states not frequently seen in the original dataset of consistently successful trials.
-Hyperparameter Analysis
+Ablation on Finetuning with Expert-Assisted Recovery Data.
+We evaluate two training configurations.
+The Original variant uses only fully successful demonstrations, while the Expert-Assisted variant is finetuned on additional trajectories where failures are corrected by an expert policy.
+As shown in Table 7, expert assisted finetuning improves performance across all difficulty levels, raising the average success rate from 82.0% to 85.6% and yielding substantial gains on Medium and Hard tasks.
+These results indicate that exposure to partial-to-successful executions enhances robustness by covering recovery scenarios absent in the original dataset.
+Overall, expert assisted augmentation provides a simple and effective boost to long-horizon reliability.
+Table 7: Comparison of task success rates (%) with and without expert assisted data augmentation across four difficulty levels in Meta-World.
+Original denotes the model trained only on the standard dataset.
+Expert-Assisted denotes the model additionally trained on expert-corrected rollouts.
+Easy Medium Hard Very Hard Average
+Original 94.8% 63.3% 21.3% 78.7% 82.0%
+Expert-Assisted 96.5% 74.9% 38.7% 84.0% 85.6%
+4.5 Error Type Analysis
+Human evaluation on MetaWorld reveals that 65% of the generated videos are correct, with errors mainly attributed to fluency issues (9.1%), robot inconsistency (11.0%), and object inconsistency (14.5%).
+For action execution, 85.5% of the predictions are correct, with the remaining errors caused by physics violations (9.1%) and spatial mistakes (5.5%).
+The low spatial error rate indicates that the spatial planning module and spatial-conditioned policy effectively capture task geometry, and most failures arise from challenging contact dynamics rather than misinterpreted spatial relations.
+4.6 Robustness to Occlusions via Masking
+Figure 6: Video generation results with partial occlusion by random masking. Despite masked regions (left), the model generates coherent future frames (right) and maintains correct task execution.
+To evaluate the robustness of our video generation module under partial visual occlusions, we introduce a masking mechanism before feeding the current frame into the diffusion model.
+Specifically, we randomly mask a portion of the image input, simulating occlusions such as those caused by environmental factors or sensor noise.
+As illustrated in Figure 6, although significant areas of the current frame are masked, the generated frames remain visually consistent and the downstream agent successfully completes the manipulation tasks.
+This indicates that the model has learned a robust representation of both visual and temporal context, and can infer plausible continuations even when critical input regions are missing.
+4.7 Hyperparameter Analysis
 Effect of Diffusion Noise Seed in Video Generation.
@@
 Setting the goal frame as a random future frame within the next $k=20$ timesteps yields over 10% higher average success rates compared to restricting it to a narrower window between 17–20 timesteps ahead, indicating that allowing a broader temporal horizon improves generalization and task adaptability.
-Second, regressing the noise $\epsilon$ achieves better performance than directly predicting the denoised action $\hat{a}_{0}$, which contrasts with some prior findings (Chi et al.
-2023).
+Second, regressing the noise $\epsilon$ achieves better performance than directly predicting the denoised action $\hat{a}_{0}$, which contrasts with some prior findings [diffusionpolicy].
 Lastly, while direct MLP-based action regression converges faster, the diffusion-based denoising model demonstrates superior performance and robustness across tasks.
-Conclusion
+5 Conclusion
 We present SP, a spatial-aware visuomotor framework that integrates explicit spatial modeling and reasoning into robot manipulation.
 Central to our approach is a spatial plan table that encodes atomic actions, directions, and distances, enabling structured spatial guidance throughout the pipeline.
-Our polciy generates spatially coherent visual trajectories by conditioning a video diffusion model on spatial plan table, , which are then used to guide action prediction based on temporally aligned spatial coordinates.
+Our policy generates spatially coherent visual trajectories by conditioning a video diffusion model on a spatial plan table, which are then used to guide action prediction based on temporally aligned spatial coordinates.
 A dual-stage replanning mechanism enhances spatial reasoning by refining the spatial plan table and regenerating video and actions upon inconsistency, ensuring closed-loop correction.
 Experiments on MetaWorld show that SP improves success rates by over 33% compared to prior methods, validating the effectiveness of spatial modeling and reasoning.
-Appendix A Inference Pipeline
+[…]
+Supplementary Material
+A.1 Inference Pipeline
 As shown in alg.
@@
 until $\texttt{VLMCheck}(\mathbf{V}_{1:T})$ succeeds;
-$\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\quad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
+$\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\qquad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
 for $t=1$ to $H$ do
@@
     until $\texttt{VLMCheck}(\mathbf{V}_{1:T})$ succeeds;
-    $\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\quad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
+    $\mathbf{z}_{\text{coord}}\leftarrow\texttt{EncodeCoord}(\mathbf{p}_{\text{ee}},\mathbf{p}_{\text{obj}}),\qquad\mathcal{A}_{1:H}\leftarrow\pi(\mathbf{V}_{1:T},\mathbf{z}_{\text{coord}})$ ;
 $\mathcal{S}\leftarrow\mathcal{E}.\texttt{CheckSuccess}()$ ;
@@
 Algorithm 1 Inference Pipeline
-Appendix B Training Details: Video Generation and Action Prediction
-Video Generation
+A.2 Training Details: Video Generation and Action Prediction
+A.2.1 Video Generation
 Training Settings.
 All models are trained on $128\times 128$ RGB frames sampled as 8-frame clips.
-We use the AdamW optimizer (Loshchilov and Hutter 2019) with an initial learning rate of $1\times 10^{-4}$ and a cosine learning rate schedule.
+We use the AdamW optimizer [adam] with an initial learning rate of $1\times 10^{-4}$ and a cosine learning rate schedule.
 Training is conducted for 100,000 steps, unless otherwise specified.
@@
 Key differences and training datasets are summarized in Table 8.
+Table 8: Comparison of video generation model variants highlighting key architectural or training differences.
 Version Description
@@
 Local Conditioning on Task & Subplan Introduces cross-attention to fuse global and local plan features for fine-grained temporal conditioning.
-Table 6: Comparison of video generation model variants highlighting key architectural or training differences.
-Action Prediction
+A.2.2 Action Prediction
 Training Settings.
-Unlike the original Diffusion Policy (Chi et al.
-2023), which conditions on past observations and actions, our model relies solely on the current observation, future observation, and spatial context.
+Unlike the original Diffusion Policy [diffusionpolicy], which conditions on past observations and actions, our model relies solely on the current observation, future observation, and spatial context.
 Specifically, instead of using a standard 3-channel RGB input, we concatenate the current frame and a sampled goal frame to form a 6-channel input image of resolution $128\times 128$, along with spatial coordinates.
 Given horizon $H=4$, the model predicts $H$ future actions from a single observation step ($T=1$), conditioned on the current frame, spatial coordinates, and a goal frame sampled within the next $K=20$ frames.
-We adopt a ResNet-50 visual encoder (He et al.
-2016) pretrained on ImageNet (Deng et al.
-2009), which remains trainable during training.
+We adopt a ResNet-50 visual encoder [resnet] pretrained on ImageNet [imagenet], which remains trainable during training.
 The spatial coordinates are encoded via a MLP, which is also trained end-to-end.
@@
 Conditioning on spatial plans significantly enhances semantic alignment and spatial fidelity, leading to more coherent and goal-directed behaviors.
-Appendix C Tasks in Meta-World
-In the Meta-World benchmark (Yu et al.
-2020), we evaluated the performance on 11 representative robotic manipulation tasks shown in Figure 5, which are designed to assess various manipulation skills and generalization abilities of reinforcement learning algorithms.
+A.3 Tasks in Meta-World and iThor
+Meta-World.
+In the Meta-World benchmark [metaworld], we evaluated the performance on 11 representative robotic manipulation tasks shown in Figure 8, which are designed to assess various manipulation skills and generalization abilities of reinforcement learning algorithms.
 These tasks are categorized by their difficulty levels as follows:
@@
 These tasks cover a range of manipulation skills, including pushing, pulling, grasping, placing, pressing, and turning, and they involve different objects and goals, providing a comprehensive evaluation of the agent’s ability to learn and generalize across various manipulation tasks.
-Figure 5: Examples of 11 tasks in meta-world
-Appendix D Dataset Construction
-We construct our dataset from expert demonstrations collected in the Meta-World (Yu et al.
-2020) simulator.
+iThor.
+To assess generalization in visually diverse, embodied environments, we further evaluate on 12 navigation-style manipulation tasks in iTHOR, spanning four rooms: Kitchen, Living Room, Bedroom, and Bathroom.
+Tasks include locating or approaching objects under varying viewpoints and scene layouts, such as Blinds, Bread, DeskLamp, Laptop, Mirror, Painting, Pillow, SoapBar, Spatula, Television, Toaster, and ToiletPaper.
+The full set of 4 rooms and 12 tasks is shown in Figure 9.
+Together, these two suites provide broad coverage of manipulation and navigation challenges, enabling a comprehensive evaluation of visuomotor generalization across simulated robotic arms and interactive 3D household environments.
+Figure 8: Generated video plans on the $1$ Meta-World tasks.
+Figure 9: Generated video plans across 12 tasks in four iTHOR room types (Kitchen, Living Room, Bedroom, Bathroom).
+A.4 Dataset Construction
+We construct our dataset from expert demonstrations collected in the Meta-World [metaworld] simulator.
 For each task, successful execution trajectories are recorded and stored as NPZ and PKL files.
@@
 The final dataset consists of (1) sampled RGB/mask/depth frames, (2) frame-wise 2D/3D positions of grippers and targets, (3) their pairwise distances, and (4) spatial plan table in textual form.
-Appendix E VLM Prompt Design
+A.5 VLM Prompt Design
 Figure 10: Detailed Design of ours VLM Prompt
@@
   2. 2.
-    Spatial plan table generation: As described in Section  D, we format spatial plan table into structured multi-step plans of the form:
-    1Plan:
+    Spatial plan table generation: As described in Section  A.4, we format spatial plan table into structured multi-step plans of the form:
+    Plan:
     1. move [-1, 0, 0] [0.19]
@@
 Instead, spatial plan table with explicit roles, response formats, and spatial semantics lead to significantly more stable and interpretable VLM outputs, which are essential for real-time feedback in visual policy generation.
-Appendix F Frame Matching Mechanism
+A.6 Frame Matching Mechanism
 To ensure precise alignment between visual goals and execution states, we introduce a similarity-based frame matching mechanism.
@@
 $\begin{split}\text{Sim}(\mathbf{I}_{\text{cur}},\hat{\mathbf{I}}_{g})=&\ w_{\text{geo}}\cdot\text{Sim}_{\text{geo}}+w_{\text{pos}}\cdot\text{Sim}_{\text{pos}}\\ &+w_{\text{ssim}}\cdot\text{SSIM}(\mathbf{I}_{\text{cur}},\hat{\mathbf{I}}_{g})+w_{\text{flow}}\cdot\text{Sim}_{\text{flow}},\end{split}$ (8)
-where $\text{Sim}_{\text{geo}}$ is the similarity based on contour and shape features (Lowe 2004), $\text{Sim}_{\text{pos}}$ encodes block-wise intensity statistics and center-of-mass shifts, SSIM is the structural similarity index between grayscale frames (Wang et al.
-2004), and $\text{Sim}_{\text{flow}}$ measures the inverse magnitude of optical flow (Farnebäck 2003).
+where $\text{Sim}_{\text{geo}}$ is the similarity based on contour and shape features [lowe2004distinctive], $\text{Sim}_{\text{pos}}$ encodes block-wise intensity statistics and center-of-mass shifts, SSIM is the structural similarity index between grayscale frames [wang2004image], and $\text{Sim}_{\text{flow}}$ measures the inverse magnitude of optical flow [farneback2003two].
 The weights are set to $w_{\text{geo}}{=}0.35$, $w_{\text{pos}}{=}0.35$, $w_{\text{ssim}}{=}0.2$, and $w_{\text{flow}}{=}0.1$.
@@
 To prevent deadlocks when similarity fails to exceed the threshold, a hard switch is enforced every $T_{\text{max}}{=}28$ steps.
-Appendix G Robustness to Occlusions via Masking
-Figure 7: Video generation results with partial occlusion by random masking. Despite masked regions (left), the model generates coherent future frames (right) and maintains correct task execution.
-To evaluate the robustness of our video generation model under partial visual occlusions, we introduce a masking mechanism before feeding the current frame into the diffusion model.
-Specifically, we randomly mask a portion of the image input, simulating occlusions such as those caused by environmental factors or sensor noise.
-As illustrated in Figure 7, although significant areas of the current frame are masked, the generated frames remain visually consistent and the downstream agent successfully completes the manipulation tasks.
-This indicates that the model has learned a robust representation of both visual and temporal context, and can infer plausible continuations even when critical input regions are missing.